# LICO: Large Language Models for In-Context Molecular Optimization

### v1

@@ 5.1 PMO Benchmark
Task GP BO Graph GA LICO REINVENT TNP
albuterol_similarity ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.636\pm 0.106}}$ $0.583\pm 0.065$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.656\pm 0.125}}$ $0.496\pm 0.020$ $0.550\pm 0.034$
amlodipine_mpo ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.519\pm 0.014}}$ $0.501\pm 0.016$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.541\pm 0.026}}$ $0.472\pm 0.008$ $0.491\pm 0.014$
celecoxib_rediscovery $0.411\pm 0.046$ $0.424\pm 0.049$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.447\pm 0.073}}$ $0.370\pm 0.029$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.429\pm 0.048}}$
deco_hop ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.593\pm 0.013}}$ $0.581\pm 0.006$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.596\pm 0.010}}$ $0.572\pm 0.006$ $0.586\pm 0.003$
drd2 ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.857\pm 0.080}}$ $0.833\pm 0.065$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.859\pm 0.066}}$ $0.775\pm 0.086$ $0.758\pm 0.066$
fexofenadine_mpo ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.707\pm 0.021}}$ $0.666\pm 0.009$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.700\pm 0.023}}$ $0.650\pm 0.007$ $0.680\pm 0.015$
isomers_c7h8n2o2 $0.545\pm 0.158$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.735\pm 0.112}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.779\pm 0.099}}$ $0.725\pm 0.064$ $0.694\pm 0.123$
isomers_c9h10n2o2pf2cl $0.599\pm 0.059$ $0.630\pm 0.086$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.672\pm 0.075}}$ $0.630\pm 0.032$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.635\pm 0.071}}$
median1 ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.213\pm 0.020}}$ $0.208\pm 0.015$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.217\pm 0.019}}$ $0.205\pm 0.012$ $0.210\pm 0.012$
median2 ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.203\pm 0.009}}$ $0.181\pm 0.009$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.193\pm 0.009}}$ $0.188\pm 0.010$ $0.186\pm 0.013$
mestranol_similarity ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.427\pm 0.025}}$ $0.362\pm 0.017$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.423\pm 0.016}}$ $0.379\pm 0.026$ $0.368\pm 0.005$
osimertinib_mpo ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.766\pm 0.006}}$ $0.751\pm 0.005$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.759\pm 0.008}}$ $0.737\pm 0.007$ $0.752\pm 0.006$
perindopril_mpo ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.458\pm 0.019}}$ $0.435\pm 0.016$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.473\pm 0.009}}$ $0.404\pm 0.009$ $0.433\pm 0.010$
qed $0.912\pm 0.010$ $0.914\pm 0.007$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.925\pm 0.005}}$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.921\pm 0.002}}$ $0.917\pm 0.002$
ranolazine_mpo ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.701\pm 0.023}}$ $0.620\pm 0.014$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.687\pm 0.029}}$ $0.574\pm 0.044$ $0.613\pm 0.033$
scaffold_hop ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.478\pm 0.009}}$ $0.461\pm 0.008$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.480\pm 0.008}}$ $0.447\pm 0.010$ $0.469\pm 0.015$
sitagliptin_mpo $0.232\pm 0.083$ $0.229\pm 0.053$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.315\pm 0.097}}$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.261\pm 0.026}}$ $0.221\pm 0.034$
thiothixene_rediscovery ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.351\pm 0.039}}$ $0.322\pm 0.023$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.343\pm 0.035}}$ $0.311\pm 0.021$ $0.307\pm 0.034$
troglitazone_rediscovery ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.313\pm 0.018}}$ $0.267\pm 0.015$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.292\pm 0.028}}$ $0.246\pm 0.009$ $0.266\pm 0.009$
valsartan_smarts $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
zaleplon_mpo $0.392\pm 0.034$ $0.374\pm 0.024$ ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.404\pm 0.022}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.406\pm 0.017}}$ $0.377\pm 0.018$
Sum of scores ($\uparrow$) ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{10.313}}$ $10.076$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{10.760}}$ $9.772$ $9.944$
Mean rank ($\downarrow$) ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{2.33}}$ $3.62$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{1.48}}$ $3.86$ $3.65$
Results Table 1 summarizes the performance of the $5$ considered methods across $21$ optimization tasks. Overall, LICO is the leading method in this benchmark, achieving the highest aggregated score and the lowest mean rank. Specifically, LICO achieves the best performance in $12/21$ tasks and the second-best performance in the remaining $8/21$ tasks. It is important to note that LICO achieves this impressive result without being explicitly trained on data from downstream objectives. This shows the effectiveness of semi-synthetic training in enabling generalization to a broad range of functions via in-context prompting. TNP performs poorly in this benchmark, despite sharing a similar architecture with LICO. This performance gap highlights the importance of the pattern-matching capabilities LLMs acquire through extensive pretraining, which are crucial for adapting to new domains.

### v2

@@ 5.1 PMO Benchmark
Task GP BO Graph GA REINVENT LICO Genetic GFN Augmented Memory MOLLEO
albuterol_similarity $0.636\pm 0.106$ $0.583\pm 0.065$ $0.496\pm 0.020$ $0.656\pm 0.125$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.664\pm 0.054}$} $0.557\pm 0.048$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.886\pm 0.023}$}
amlodipine_mpo $0.519\pm 0.014$ $0.501\pm 0.016$ $0.472\pm 0.008$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.541\pm 0.026}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.534\pm 0.019}$} $0.489\pm 0.009$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.637\pm 0.023}$}
celecoxib_rediscovery $0.411\pm 0.046$ $0.424\pm 0.049$ $0.370\pm 0.029$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.447\pm 0.073}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.447\pm 0.028}$} $0.385\pm 0.027$ $0.402\pm 0.003$
deco_hop $0.593\pm 0.013$ $0.581\pm 0.006$ $0.572\pm 0.006$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.596\pm 0.010}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.604\pm 0.017}$} $0.579\pm 0.010$ $0.588\pm 0.007$
drd2 $0.857\pm 0.080$ $0.833\pm 0.065$ $0.775\pm 0.086$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.859\pm 0.066}$} $0.809\pm 0.045$ $0.795\pm 0.024$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.910\pm 0.017}$}
fexofenadine_mpo $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.707\pm 0.021}$} $0.666\pm 0.009$ $0.650\pm 0.007$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.700\pm 0.023}$} $0.682\pm 0.021$ $0.679\pm 0.021$ $0.674\pm 0.002$
gsk3b $0.611\pm 0.059$ $0.523\pm 0.047$ $0.589\pm 0.063$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.617\pm 0.063}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.637\pm 0.018}$} $0.539\pm 0.097$ $0.397\pm 0.013$
isomers_c7h8n2o2 $0.545\pm 0.158$ $0.735\pm 0.112$ $0.725\pm 0.064$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.779\pm 0.099}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.738\pm 0.039}$} $0.661\pm 0.039$ $0.737\pm 0.043$
isomers_c9h10n2o2pf2cl $0.599\pm 0.059$ $0.630\pm 0.086$ $0.630\pm 0.032$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.672\pm 0.075}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.656\pm 0.075}$} $0.596\pm 0.066$ $0.635\pm 0.017$
jnk3 $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.346\pm 0.067}$} $0.301\pm 0.071$ $0.315\pm 0.042$ $0.336\pm 0.051$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.409\pm 0.165}$} $0.294\pm 0.110$ $0.186\pm 0.076$
median1 $0.213\pm 0.020$ $0.208\pm 0.015$ $0.205\pm 0.012$ $0.217\pm 0.019$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.219\pm 0.008}$} $0.219\pm 0.014$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.236\pm 0.021}$}
median2 $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.203\pm 0.009}$} $0.181\pm 0.009$ $0.188\pm 0.010$ $0.193\pm 0.009$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.204\pm 0.011}$} $0.184\pm 0.010$ $0.191\pm 0.009$
mestranol_similarity $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.427\pm 0.025}$} $0.362\pm 0.017$ $0.379\pm 0.026$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.423\pm 0.016}$} $0.414\pm 0.022$ $0.393\pm 0.021$ $0.399\pm 0.020$
osimertinib_mpo $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.766\pm 0.006}$} $0.751\pm 0.005$ $0.737\pm 0.007$ $0.759\pm 0.008$ $0.763\pm 0.008$ $0.761\pm 0.006$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.779\pm 0.006}$}
perindopril_mpo $0.458\pm 0.019$ $35\pm 0.016 0.404$$04\pm 0.009$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.473\pm 0.009}$} $0.462\pm 0.033$ $0.422\pm 0.013$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.655\pm 0.054}$}
qed $0.912\pm 0.010$ $0.914\pm 0.007$ $0.921\pm 0.002$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.925\pm 0.005}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.928\pm 0.002}$} $0.923\pm 0.002$ $0.919\pm 0.006$
ranolazine_mpo $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.701\pm 0.023}$} $0.620\pm 0.014$ $0.574\pm 0.044$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.687\pm 0.029}$} $0.623\pm 0.022$ $0.614\pm 0.033$ $0.640\pm 0.000$
scaffold_hop $0.478\pm 0.009$ $0.461\pm 0.008$ $0.447\pm 0.010$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.480\pm 0.008}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.485\pm 0.015}$} $0.460\pm 0.010$ $0.473\pm 0.000$
sitagliptin_mpo $0.232\pm 0.083$ $0.229\pm 0.053$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.261\pm 0.026}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.315\pm 0.097}$} $0.227\pm 0.041$ $0.245\pm 0.030$ $0.193\pm 0.073$
thiothixene_rediscovery $0.351\pm 0.039$ $0.322\pm 0.023$ $0.311\pm 0.021$ $0.343\pm 0.035$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.377\pm 0.015}$} $0.336\pm 0.033$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.416\pm 0.075}$}
troglitazone_rediscovery $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.313\pm 0.018}$} $0.267\pm 0.015$ $0.246\pm 0.009$ $0.292\pm 0.028$ $0.277\pm 0.015$ $0.262\pm 0.012$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.302\pm 0.022}$}
valsartan_smarts $\pm 0.000 0.000±0.000$$\pm 0.000$ $\pm 0.000 0.000±0.000$$\pm 0.000$ $\pm 0.000 0.000±0.000$$\pm 0.000$ $0.000\pm 0.000$
zaleplon_mpo $0.392\pm 0.034$ $0.374\pm 0.024$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.406\pm 0.017}$} $0.404\pm 0.022$ $0.400\pm 0.014$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.415\pm 0.013}$} $0.392\pm 0.003$
Sum of scores ($\uparrow$) $11.27$ $10.90$ $10.68$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{11.71}$} $11.56$ $10.81$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{11.65}$}
Table 2: The performance of LICO and the baselines on $23$ optimization tasks in PMO. A higher score is better. We report the mean and stddev of scores averaged over 55 random seeds. We use blue and violet to denote the best and second-best method for each task.
Task GP BO Graph GA REINVENT LICO Genetic GFN Augmented Memory MOLLEO
albuterol_similarity $0.898\pm 0.014$ $0.838\pm 0.016$ $0.882\pm 0.006$ $0.885\pm 0.019$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.941\pm 0.021}$} $0.913\pm 0.009$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.936\pm 0.016}$}
amlodipine_mpo $0.583\pm 0.044$ $0.661\pm 0.020$ $0.635\pm 0.035$ $0.679\pm 0.027$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.709\pm 0.027}$} $0.691\pm 0.047$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.801\pm 0.028}$}
celecoxib_rediscovery $0.723\pm 0.053$ $0.630\pm 0.097$ $0.713\pm 0.067$ $0.664\pm 0.122$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.784\pm 0.032}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.796\pm 0.008}$} $0.459\pm 0.080$
deco_hop $0.629\pm 0.018$ $0.619\pm 0.004$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.666\pm 0.044}$} $0.619\pm 0.015$ $0.653\pm 0.028$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.658\pm 0.024}$} $0.648\pm 0.099$
drd2 $0.923\pm 0.017$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.964\pm 0.012}$} $0.945\pm 0.007$ $0.928\pm 0.018$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.963\pm 0.006}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.963\pm 0.006}$} $0.962\pm 0.013$
fexofenadine_mpo $0.722\pm 0.005$ $0.760\pm 0.011$ $0.784\pm 0.006$ $0.772\pm 0.023$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.793\pm 0.009}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.859\pm 0.009}$} $0.776\pm 0.019$
gsk3b $0.851\pm 0.041$ $0.788\pm 0.070$ $0.865\pm 0.043$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.876\pm 0.045}$} $0.861\pm 0.022$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.881\pm 0.021}$} $0.865\pm 0.037$
isomers_c7h8n2o2 $0.680\pm 0.117$ $0.862\pm 0.065$ $0.852\pm 0.036$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.939\pm 0.022}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.955\pm 0.007}$} $0.853\pm 0.087$ $0.915\pm 0.036$
isomers_c9h10n2o2pf2cl $0.469\pm 0.180$ $0.719\pm 0.047$ $0.642\pm 0.054$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.819\pm 0.039}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.876\pm 0.018}$} $0.736\pm 0.051$ $0.708\pm 0.093$
jnk3 $0.564\pm 0.155$ $0.553\pm 0.136$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.783\pm 0.023}$} $0.731\pm 0.037$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.759\pm 0.063}$} $0.739\pm 0.110$ $0.715\pm 0.026$
median1 $0.301\pm 0.014$ $0.294\pm 0.021$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.356\pm 0.009}$} $0.291\pm 0.016$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.355\pm 0.009}$} $0.326\pm 0.013$ $0.302\pm 0.031$
median2 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.297\pm 0.009}$} $0.273\pm 0.009$ $0.276\pm 0.008$ $0.280\pm 0.019$ $0.289\pm 0.007$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.291\pm 0.008}$} $0.206\pm 0.015$
mestranol_similarity $0.627\pm 0.089$ $0.579\pm 0.022$ $0.618\pm 0.048$ $0.614\pm 0.064$ $0.697\pm 0.035$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.750\pm 0.049}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.759\pm 0.102}$}
osimertinib_mpo $0.787\pm 0.006$ $0.831\pm 0.005$ $0.837\pm 0.009$ $0.820\pm 0.012$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.846\pm 0.008}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.855\pm 0.004}$} $0.819\pm 0.020$
perindopril_mpo $0.493\pm 0.011$ $0.538\pm 0.009$ $0.537\pm 0.016$ $0.557\pm 0.028$ $0.595\pm 0.011$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.613\pm 0.015}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.723\pm 0.018}$}
qed $0.937\pm 0.000$ $0.940\pm 0.000$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.941\pm 0.000}$} $0.936\pm 0.001$ $0.937\pm 0.000$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.942\pm 0.000}$} $0.933\pm 0.003$
ranolazine_mpo $0.735\pm 0.013$ $0.728\pm 0.012$ $0.760\pm 0.009$ $0.774\pm 0.008$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.810\pm 0.011}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.801\pm 0.006}$} $0.731\pm 0.023$
scaffold_hop $0.548\pm 0.019$ $0.517\pm 0.007$ $0.560\pm 0.019$ $0.547\pm 0.026$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.585\pm 0.041}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.567\pm 0.008}$} $0.516\pm 0.022$
sitagliptin_mpo $0.186\pm 0.055$ $433\pm 0.075 0.0$$021\pm 0.003$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.567\pm 0.034}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.577\pm 0.036}$} $0.284\pm 0.050$ $0.496\pm 0.020$
thiothixene_rediscovery $0.559\pm 0.027$ $0.479\pm 0.025$ $0.534\pm 0.013$ $0.514\pm 0.037$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.599\pm 0.073}$} $0.550\pm 0.041$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.658\pm 0.024}$}
troglitazone_rediscovery $0.410\pm 0.015$ $0.390\pm 0.016$ $0.441\pm 0.032$ $0.380\pm 0.026$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.455\pm 0.016}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.540\pm 0.048}$} $0.352\pm 0.040$
[…]
zaleplon_mpo $0.221\pm 0.072$ $0.346\pm 0.032$ $0.358\pm 0.062$ $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.515\pm 0.017}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.545\pm 0.023}$} $0.394\pm 0.026$ $0.402\pm 0.019$
Sum of scores ($\uparrow$) $13.156$ $13.751$ $14.196$ $14.708$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{15.678}$} $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{15.002}$} $14.682$
Results Table 1 summarizes the performance of the $7$ considered methods across $23$ optimization tasks in PMO-1K. Overall, LICO is the leading method in this benchmark, achieving the highest aggregated score. Specifically, LICO achieves the best or second-best performance in $14/23$ tasks. MOLLEO performs competitively with LICO on this benchmark, with a sum score of $11.65$. However, we note that MOLLEO has significant advantages over LICO and other methods. MOLLEO used BioT5 to generate valid molecules, which has been finetuned extensively on molecules, protein, and molecule-related text data, while LICO leveraged a general LLM like Llama. Moreover, MOLLEO prompts the LLM with a detailed textual description of the task such as "Your job is to generate a SELFIES molecule that looks more like a drug", which possibly has data contamination issues, since the finetuning data may have included similar tasks. On the other hand, we use LICO as a black-box surrogate model, where the model makes predictions based purely on in-context learning of the mapping between molecule fingerprints and their corresponding scores.
On the original PMO setting, Table 2 shows the competitive performance of LICO with the two state-of-the-art methods Genetic GFN and Augmented Memory. It is important to note that other methods have significant advantages over LICO, since both Genetic GFN and Augmented Memory update their models from real data during optimization, whereas LICO performs in-context learning without being explicitly trained on data from downstream objectives. This impressive result shows the effectiveness of semi-synthetic training in enabling generalization to a broad range of functions via in-context prompting.